# Tailoring Polarization in WSe Quantum Emitters through Deterministic Strain Engineering

### 16 Feb 2024 v1

@@ Abstract
A quantum emitter in a transition metal dichalcogenide has recently emerged as a promising platform for generating single photons for optical quantum information processing. In this work, we present an approach for deterministically controlling the polarization of fabricated quantum emitters in a tungsten diselenide (WSe) monolayer. We employ unconventional nanopillar geometries with high apex-to-base ratios and sharp vertices to induce a controlled directional strain in the monolayer, and we report on fabricated WSe emitters producing single photons with a high degree of polarization and high purity (). Our work paves the way for the deterministic integration of TMD-based quantum emitters in quantum photonic architecture.
@@ I Introduction
Within photonic quantum information technologies, a key component is the source of single photons Gregersen et al. (2017) used to encode the quantum bits. The ideal single-photon source (SPS) should feature deterministic emission of single indistinguishable photons with near-unity collection efficiency and high purity. The spontaneous parametric down-conversion process is a straightforward method Kwiat et al. (1995) for producing highly indistinguishable photons and has been the workhorse of the community for several decades; however, its probabilistic nature results in a significant trade-off between photon purity and efficiency. As an alternative, the two-level system in a semiconductor host material Aharonovich et al. (2016) has recently emerged as an attractive platform for the deterministic generation of single photons using the spontaneous emission process. By placing a semiconductor quantum dot (QD) in an optical cavity Heindel et al. (2023) and by exploiting cavity quantum electrodynamics (cQED) to control the light emission, emission of pure highly indistinguishable photons with an efficiency of 0.6 has been demonstrated Tomm et al. (2021); Wang et al. (2019). However, the fabrication of high-quality semiconductor QDs requires expensive unreliable molecular beam epitaxy growth methods representing a major drawback in the development of QD-based sources.
@@
As an alternative, quantum emitters in transition metal dichalcogenides (TMDs) are attractive due to their availability, their ease of integration into nanophotonic structures, and the versatility in engineering their photonic characteristics Montblanch et al. (2023). While the direct bandgap structure Le et al. (2015); Montblanch et al. (2023) of the pristine monolayer form allows TMDs to act as an efficient classical light source, the emission of pure single photons requires additional engineering. The microscopic origin of single-photon emission in TMDs is believed to be localized excitonic states Aharonovich et al. (2016); Gao et al. (2023); Iff et al. (2019); Linhart et al. (2019); Parto et al. (2021); Branny et al. (2017) appearing either due to defects in the crystal lattice or due to the application of strain. Thus the implementation of single-photon emitters (SPEs) into TMDs has been pursued by using deterministic defect fabrication Parto et al. (2021); Xu et al. (2023); Wang et al. (2022) as well as by introducing localized strain in TMD monolayers Kumar et al. (2015); Kern et al. (2016). While stress can be applied directly to the flat TMD lattice Iff et al. (2019); So et al. (2021); Xu et al. (2023); Yu et al. (2021), the strain is typically introduced by placing the TMD layer on arrays of nanopillars Parto et al. (2021); Iff et al. (2021); Branny et al. (2017); Palacios-Berraquero et al. (2017); Kim et al. (2019); Zhao et al. (2021); Luo et al. (2018); Azzam et al. (2023); Kern et al. (2016) leading to strain concentration at the contact point.
@@
Similar to QDs, efficient single-photon emission can be obtained from an SPE in a TMD by placing it inside an optical cavity and exploiting cQED. A WSe emitter was recently placed inside an open cavity geometry Drawer et al. (2023) leading to the demonstration of an efficiency of 0.65 as well as 0.02 indistinguishability. The cQED effect relies on a matching of polarizations of the dipole of the emitter and the electric field of the optical cavity mode. For the strain-induced SPE, a deformation strain potential, a nanowrinkle, serving as a macroscopic host for the SPE, is produced by the contact of the WSe monolayer with the nanopillar. The polarization of the SPE dipole is then aligned So et al. (2021); Azzam et al. (2023); Wang et al. (2021) with the orientation of the wrinkle. Control of the lateral dipole orientation is not needed in a rotationally symmetric geometry Drawer et al. (2023) with two degenerate orthogonally polarized optical cavity modes, where the dipole can couple into either mode. However, the general case requires control of the SPE dipole polarization, and thus the wrinkle orientation, to efficiently couple light into, for example, the polarized TE mode of a planar ridge waveguide as needed for on-chip integrated quantum photonics Wang et al. (2020).
In this work, we propose a simple and efficient method for fabricating orientation-controlled wrinkles using strain-engineered monolayer WSe, and we demonstrate light emission from SPEs with high purity () and a high degree of polarization (99 4%). Our method is based on nanostructures tailored to deterministically induce strain in a specific direction. Moreover, we artificially introduce lattice defects with electron beam irradiation to further enhance the emission process. This work paves the way for the controlled integration of TMD-based quantum emitters into photonic systems.
II Results and Discussions
II.1 Engineering the WSe quantum emitters via cylindrical nanopillars
The material used in this study, WSe, has an emission wavelength range (750-800 nm) compatible with low-loss free-space channels and with the wavelength of Rb-87-D2 for integration with quantum memories Drawer et al. (2023); Simon et al. (2010).
We initially fabricated single-photon emitters on conventional Parto et al. (2021); Palacios-Berraquero et al. (2017) rotationally symmetric nanopillars. Pristine monolayer flakes of WSe are transferred on top of an array of cylindrical nanopillars using the dry transfer technique Castellanos-Gomez et al. (2014). The pillar dimensions were set to 150 nm in height and 0.5 µm in diameter. To adapt to the shape of the pillar, the flakes undergo mechanical strain, deform, and create nanowrinkles. The detailed fabrication process is discussed in the Methods section as well as in Figure S1 and Supporting Note 1. Bright field (BF) and photoluminescence (PL, T = 4 K) images of the fabricated sample with cylindrical pillars are shown in Figures 1a,b, respectively. In Figure 1a, a white outline highlights the monolayer, while the pillars are identified through their contrast difference, appearing as dark spots. As can be seen in Figure 1b, a brighter PL emission signal surrounding the pillar was observed, in comparison with the planar region of WSe. This PL increase could be attributed to the efficient migration of excitons toward lower-energy states in the strained areas. Interestingly, the brighter PL emission signal seems to be located around the pillar. This suggests that a stronger strain is localized around the pillar, not on top, which is in good agreement with the literature Branny et al. (2017). However, it is unclear from the PL image how the nanopillar induces a local strain in the monolayer.
@@ II.1 Engineering the WSe quantum emitters via cylindrical nanopillars
To provide microscopic details of the deformation around the pillar, we performed the scanning electron microscope (SEM) imaging of the fabricated sample, presented in Figure 1c. The obtained SEM image reveals the formation of four straight nanowrinkles around the pillar.
@@
It is generally accepted that these nano-wrinkles host quantum emitters due to local strain, and their orientation is what dictates the polarization profile of the emitted single photons from the resulting emitters So et al. (2021); Azzam et al. (2023); Wang et al. (2021).
@@
Due to the rotational symmetry of the nanopillar, which applies strain uniformly, the final number and orientation of these nanowrinkles are random, rendering the results of each sample irreproducible. Moreover, their close proximity and difficult spatial localization of nanowrinkles around the nanopillar could potentially affect further optical characterizations, e.g., purity.
@@
To investigate the impact of closely formed nanowrinkles in the presence of a single pillar on the optical properties of the sample, we measured PL spectra (T = 4 K) from five different spots around the pillar, highlighted with white crosses (S1-S5) and present them in Figure 1d. A broad emission signal is observed from the planar monolayer region (S1) due to the delocalized excitons, whereas the collected spectra taken from different spots around the nanowrinkles (S2-S5) exhibit new discrete and narrow emission lines with higher intensities ranging from 730 - 810 nm. These narrow emission lines are attributed to the new localized excitons hosted in the confined systems produced by the strain and defects in the mono- and a few layers WSe Montblanch et al. (2023); Dang et al. (2020); Xu et al. (2023); Linhart et al. (2019); Parto et al. (2021).
The broad emission, similar to the planar monolayer, was also observed from the nanowrinkles (S2-S5) in the spectral range between 720-770 nm due to the collection spot area (larger than the emitter region), which leads to partial collection from the planar region. This makes isolating individual emission lines associated with potential single-photon emission from the broad multiphoton emission background relatively challenging, even with narrow optical bandpass filters. However, since these narrow lines appear relatively intense, it still might be possible to perform further characterization. Moreover, due to the spatial proximity of the nanowrinkles, the PL spectra exhibit overlapping emission lines, as we capture light emitted from multiple emitters simultaneously. Furthermore, these features might share their in-plane polarization axis if the nanowrinkles hosting them have a parallel orientation (S2-S4, S3-S5), making their spectral isolation practically impossible. Nevertheless, we found several intense individual emission lines in longer wavelengths above 780 nm, which could be easily isolated using a long-pass optical filter. One crucial characteristic of quantum emitters, highly relevant to numerous quantum photonic applications, is the degree of purity, indicating whether they emit only one photon at a time. This single-photon purity is often assessed through the value. To assess the purity of these new emission lines in Figure 1d, we performed the second-order autocorrelation () measurement from a representative emission line ( nm) under continuous-wave (CW) excitation at 455 nm via a Hanbury–Brown–Twiss (HBT) setup. In this setup, the emitted photons are coupled to a 50:50 fiber beam splitter, whose output signals are detected by superconducting nanowire single-photon detectors (SNSPD). The emission line at 807 nm was spectrally isolated using a 750 nm long-pass filter. Figure 1e shows the obtained correlation histogram, exhibiting a clear signature of photon antibunching. The extracted value of is below the threshold, indicating single-photon emission behavior. Despite the apparent successful fabrication of SPEs with cylindrical pillars, the challenges posed by the irreproducibility of results and the lack of deterministic control of polarization in TMD-based emitters underscore the need for an effective strain-engineering approach to potentially overcome these limitations. In the following section, we discuss a unique method of employing novel nanostructure designs to induce directional strain in the WSe monolayer in a deterministic manner for controlling the in-plane polarization and spatial localization of the WSe emitters.
II.2 Novel nanopillar designs
@@
To address the aforementioned limitations, we propose unique nanostructure designs with tailored geometries for deterministic strain engineering to fabricate highly-polarized SPEs from monolayer WSe. The proposed pillar designs consist of polygonal structures with long vertices, sharp edges, and a relatively high apex-to-base ratio. Their shape allows us to induce the strains on the tips of the vertices and facilitate the creation of spatially isolated nanowrinkles with controlled orientation, which follows the directionality of the vertices (see Figure 2a). Controlling the direction of these one-dimensional nanowrinkles, which may host emitters, gives rise to an effective manipulation of the emitted linearly-polarized single photons So et al. (2021); Azzam et al. (2023); Wang et al. (2021). Here, we explore three different pillar geometries, namely a three-pointed star (TS), a five-pointed star (FS), and a bowtie (BT). The structures are patterned out of negative electron beam resist (HSQ), which is cured into SiO upon e-beam exposure. The thickness of the nanostructures ( nm) was identified as optimal for producing nanowrinkles during the transfer process. Further details of the e-beam lithography and the remaining fabrication process are described in the Methods section.
@@
First, we demonstrate WSe single-photon emitters with a TS pillar. Figure 2a shows a 3D schematic representation of the TS nanostructure with a WSe monolayer flake, illustrating the formation of nanowrinkles that follow the directionality of each vertex. The surface morphology of a sample hosting a WSe monolayer flake is shown in the atomic force microscope (AFM) image (see Figure 2b). The color bar accompanying the AFM image highlights the height of the TS structure while emphasizing the height difference between the fabricated nanowrinkles and the planar monolayer. We examine the morphological profiles of the fabricated nanowrinkles on this exemplary structure in Figure S3 of the SI. As shown in Figure 2b, the monolayer forms nanowrinkles mostly following the directionality of the vertices. Moreover, the nanowrinkles are separated by more than 2 µm, allowing us to address each emitter individually.
To investigate how these nanowrinkles modify the optical properties of monolayer WSe, we performed PL imaging under a 470 nm light-emitting diode (LED) excitation in a closed-cycle optical cryostat equipped with a microscope objective (60, NA = 0.82) operating at T=4 K.
The obtained PL image (T=4 K) taken from the TS sample is presented in Figure 2c, where increased PL emission is observed along the nanowrinkles. This points to the existence of single-photon emitters due to the formed directional strains in the sample.
The emission in the area of the TS structure, highlighted by a semi-transparent star, is comparable to that of the planar monolayer, confirming that it is solely produced by the nanowrinkles. More notably, the increased PL signal is not only focused on the tips of the vertices but can also be seen across the length of the nanowrinkles.
To gain further insights into the effect of the formed nanowrinkles on the polarization of emitted photons from the spatially separated nanowrinkles in the TS sample, we performed the polarization-resolved µPL measurement from the nanowrinkles using a confocal microscope setup.
The collected PL spectra from the individual nanowrinkles (P1-P3) under parallel and orthogonal polarizations with respect to the directionality of each nanowrinkle are presented in Figures 2d-f. They exhibit highly polarized emission, as the PL intensity of the orthogonal polarization is orders of magnitude lower than the parallel one. This behavior implies that the polarization orientation follows the directionality of the nanowrinkles.
To gain a better understanding, we analyzed the polarization-resolved µPL spectra taken from the individual nanowrinkles. The extracted PL intensity as a function of the in-plane polarization exhibits highly linear polarization behavior, as shown in Figure 2g-i. To quantify the degree of linear polarization (DOLP) of the emitted photons, we use a sinusoidal fitting function and extract the DOLP using the maximum and minimum PL intensities as , with the calculated values of 99 ± 4%, 82 ± 14%, and 93 ± 3%, respectively. The strong linear polarization is attributed to the nanowrinkles, which give rise to a localized one-dimensional strain potential, consequently forming strongly aligned dipoles. Tripathi et al. (2018); Drawer et al. (2023) In addition, the fitting reveals that the in-plane polarization profile of each emitter follows the directionality of their respective nanowrinkles.
Based on these results, it is evident that the proposed design facilitates the successful isolation of the emitters both spatially and in terms of polarization.
More importantly, our method indeed enables reproducible and effective control of the in-plane polarization of the emitted photons by deterministically inducing directional strain in the nanowrinkle with sharp vertices.
@@
Figure 3, as in the TS case, presents a similar investigation of the single photon emission property of WSe on the FS structure. The AFM image in Figure 3 showcases the surface morphology of the FS sample and the resulting nanowrinkles. The false-shading representation is shown in the inset for better contrast. We notice that one vertex out of the five vertices does not have a visible nanowrinkle forming at its end. This is dismissed as a normal occurrence due to the intricacies of the transfer process. Moreover, we observe that in the other four vertices (P1-P4) the nanowrinkles are significantly shorter than that of the TS, which is attributed to the increased number of vertices in the FS that may obstruct the formation of longer nanowrinkles. Nevertheless, the nanowrinkles are well separated and aligned with the vertices.
@@
To further analyze the optical properties of the monolayer WSe on the FS structure, we obtain the corresponding PL image ( K) under a 470 nm LED excitation, as shown in Figure 3b. The PL signal significantly increases around the tips where the nanowrinkles are localized (P1-P4), with the tip missing a nanowrinkle showing emission comparable to the pristine monolayer. While we understand that the presence of SPEs relies on the one-dimensional potential induced via strain engineering, it becomes evident that the formation of a nanowrinkle is a prerequisite.
@@
To understand the effect of the nanowrinkles of the FS structure on the in-plane polarization of the single photons and compare it to that of the TS, we obtain polarization-resolved µPL spectra from the four highlighted spots. We extract the PL intensities and present them as polar plots in Figures 3c-f. Similarly to Figure 2, the sinusoidal fitting reveals that the emitted photons from each SPE demonstrate high DOLP, with the extracted values at 86 , 81 , 76 , and 85 , respectively. More importantly, the orientation of the in-plane polarization follows the directionality of the vertices, verifying our ability to control the polarization of the SPEs.
@@
While these structures demonstrate promising control over both the SPEs’ localization and polarization, it is worth emphasizing that the control of wrinkle directionality is not flawless. As shown in Figures 2 and 3, possible misalignments between the vertex and the nanowrinkle are possible. This behavior is attributed to the hexagonal lattice structure of the material and how its intrinsic “armchair” and “zigzag” directions Le and Rahman (2013) might affect the final directionality of the nanowrinkles. Statistical analysis of the deviation in alignment between the directionality of the vertices and the nanowrinkles of the TS and FS structures is presented in Supporting information Note S4 and Figure S4. Despite this analysis indicating that there are no overbearing misalignment issues, an upgraded design is presented in Figure 4. The BT design facilitates the formation of the nanowrinkle between the two triangles. This way, any potential misalignment between the vertices and the nanowrinkle would be minimized while isolating the SPEs from any naturally forming deformation in the surrounding area. Figure 4a shows the AFM image of an exemplary BT structure hosting the monolayer WSe flake. A clear nanowrinkle connecting the two triangles is visible.

### 20 Mar 2024 v2

@@ Abstract
Quantum emitters in transition metal dichalcogenides (TMDs) have recently emerged as a promising platform for generating single photons for optical quantum information processing. In this work, we present an approach for deterministically controlling the polarization of fabricated quantum emitters in a tungsten diselenide (WSe) monolayer. We employ novel nanopillar geometries with long and sharp tips to induce a controlled directional strain in the monolayer, and we report on fabricated WSe emitters producing single photons with a high degree of polarization and high purity (). Our work paves the way for the deterministic integration of TMD-based quantum emitters for future photonic quantum technologies.
@@ I Introduction
Within photonic quantum information technologies, a key component is the source of single photons Gregersen et al. (2017) used to encode the quantum bits. The ideal single-photon source (SPS) should feature deterministic emission of single indistinguishable photons with near-unity collection efficiency and high purity. The spontaneous parametric down-conversion process is a straightforward method Kwiat et al. (1995) for producing highly indistinguishable photons and has been the workhorse of the community for several decades; however, its probabilistic nature results in a significant trade-off between photon purity and efficiency. As an alternative, the two-level system in a semiconductor host material Aharonovich et al. (2016) has recently emerged as an attractive platform for the deterministic generation of single photons using the spontaneous emission process. By placing a semiconductor quantum dot (QD) in an optical cavity Heindel et al. (2023) and by exploiting cavity quantum electrodynamics (cQED) to control the light emission, emission of pure highly indistinguishable photons have been generated with efficiency as high as 0.6 Tomm et al. (2021); Wang et al. (2019). However, the fabrication of high-quality semiconductor QDs requires expensive, complex molecular beam epitaxy growth methods, which represent a significant drawback in developing QD-based sources.
@@
As an alternative, quantum emitters in transition metal dichalcogenides (TMDs) are attractive due to their availability, their ease of integration into nanophotonic structures, and the versatility in engineering their photonic characteristics Montblanch et al. (2023). While the direct bandgap structure Le et al. (2015); Montblanch et al. (2023) of the pristine monolayer form of TMDs allows for an efficient classical light generation, the emission of pure single photons requires additional engineering. The microscopic origin of single-photon emission in TMDs is believed to be from localized excitonic states appearing either due to defects in the crystal lattice or due to the local strainAharonovich et al. (2016); Gao et al. (2023); Iff et al. (2019); Linhart et al. (2019); Parto et al. (2021); Branny et al. (2017). Thus, the implementation of single-photon emitters (SPEs) into TMDs has been pursued by using deterministic defect fabrication Parto et al. (2021); Xu et al. (2023); Wang et al. (2022) as well as by introducing localized strain in TMD monolayers Kumar et al. (2015); Kern et al. (2016). While stress can be applied directly to the flat TMD lattice Iff et al. (2019); So et al. (2021); Xu et al. (2023); Yu et al. (2021), strain is typically introduced by placing the TMD layer on arrays of nanopillars Parto et al. (2021); Iff et al. (2021); Branny et al. (2017); Palacios-Berraquero et al. (2017); Kim et al. (2019); Zhao et al. (2021); Luo et al. (2018); Azzam et al. (2023); Kern et al. (2016) leading to strain concentration at the contact point.
@@
Similar to QDs, efficient single-photon emission can be obtained from an SPE in a TMD by placing it inside an optical cavity and exploiting cQED. A WSe emitter was recently placed inside an open cavity geometry Drawer et al. (2023) leading to the demonstration of an efficiency of 0.65 and an indistinguishability of 0.02. The cQED effect relies on a matching of polarizations of the dipole of the emitter and the electric field of the optical cavity mode. For the strain-induced SPE, a deformation strain potential—a nanowrinkle, serving as a macroscopic host for the SPE, is produced by the contact of the WSe monolayer with the nanopillar. The polarization of the SPE dipole is then aligned So et al. (2021); Azzam et al. (2023); Wang et al. (2021) with the orientation of the wrinkle. Control of the lateral dipole orientation is unnecessary in a rotationally symmetric geometry Drawer et al. (2023) with two degenerate orthogonally polarized optical cavity modes, where the dipole can couple into either mode. However, the general case requires control of the SPE dipole polarization, and thus the wrinkle orientation, to efficiently couple light into, for example, the polarized TE mode of a planar ridge waveguide as needed for on-chip integrated quantum photonics Wang et al. (2020).
In this work, we propose an efficient method for fabricating orientation-controlled nanowrinkles in monolayer WSe, and we demonstrate single photon emission with high purity () and a high degree of polarization (99 4%). This method consists of transferring flakes onto a nanostructure with long and sharp tips to form one-dimensional (1D) nanowrinkles, thereby controlling the direction of 1D strain and the polarization of the emitted photons. This work paves the way for integrating TMD-based quantum emitters into photonic systems.
II Results and Discussion
II.1 WSe quantum emitters via cylindrical nanopillars
We initially fabricate SPEs by transferring monolayer WSe flakes on an array of conventional, Parto et al. (2021); Palacios-Berraquero et al. (2017) rotationally symmetric nanopillars.
The pillar dimensions were set to 150 nm in height and 500 nm in diameter. During the transfer, the monolayer adapts to the shape of the pillar due to the application of mechanical strain and is deformed by creating nanowrinkles.
The detailed fabrication process is discussed in the Methods section and Figure S1 and Supporting Note 1. In addition to strain engineering, we introduce defects to the monolayer WSe via e-beam irradiation. Such a process is motivated in Supporting Note 2 and Figure S2 of the Supporting Information.
Figure 1a shows a bright-field (BF) image of the fabricated sample with cylindrical pillars, where the WSe monolayer is highlighted with a white outline, and the pillars are identified through their contrast difference, appearing as dark spots. Using a 470 nm light-emitting diode (LED) as an excitation source, we perform photoluminescence (PL, T = 4 K) imaging on the same sample and present it in Figure 1b. We observe increased photon emission in proximity of a nanopillar (green circle) compared to the emission in the planar region of WSe.
This PL increase could be attributed to the efficient migration of excitons toward lower-energy states in the strained areas. Interestingly, the brighter PL emission signal seems to be located around the pillar. This suggests that a stronger strain is localized around the pillar, not on top, which is in good agreement with previously reported resultsBranny et al. (2017). However, the PL image does not clarify how the nanopillar induces local strain in the monolayer.
@@ II.1 WSe quantum emitters via cylindrical nanopillars
To provide microscopic details of the deformation around the pillar, we performed scanning electron microscope (SEM) imaging of the fabricated sample, presented in Figure 1c. The obtained SEM image reveals the formation of four straight nanowrinkles around the pillar.
@@
It is generally accepted that these nanowrinkles host quantum emitters due to local strain, and their orientation is what dictates the polarization profile of the emitted single photons from the resulting emitters So et al. (2021); Azzam et al. (2023); Wang et al. (2021).
@@
Due to the rotational symmetry of the nanopillar, which applies strain uniformly, the final number and orientation of these nanowrinkles are random, rendering us unable to control these attributes. Moreover, their proximity could affect further optical characterizations, such as purity.
@@
To investigate the impact of closely formed nanowrinkles in the presence of a single pillar on the optical properties of the sample, we measured PL spectra (T = 4 K) from five different spots around the pillar, highlighted with white crosses (S1-S5) and present them in Figure 1d. A broad emission signal is observed from the planar monolayer region (S1) due to the delocalized exciton and trion states, whereas the collected spectra taken from different spots around the nanowrinkles (S2-S5) exhibit new discrete and narrow emission lines with higher intensities ranging from 730 to 810 nm Palacios-Berraquero et al. (2017). These narrow emission lines are attributed to the new localized excitons hosted in the confined systems produced by the strain and defects in the mono- and few-layer WSe Montblanch et al. (2023); Dang et al. (2020); Xu et al. (2023); Linhart et al. (2019); Parto et al. (2021); Kern et al. (2016); Iff et al. (2018); Errando-Herranz et al. (2021).
Due to the collection area being larger than the emitter region, we partially collect light from the planar monolayer region, resulting in the broad emission being observed in the spectra from the nanowrinkles (S2-S5) ranging between 720-770 nm. This makes isolating individual emission lines associated with potential single-photon emission from the broad multiphoton emission background relatively challenging, even with narrow optical band-pass filters. Moreover, due to the spatial proximity of the nanowrinkles, the PL spectra exhibit overlapping emission lines as we capture light emitted from multiple emitters simultaneously. These emission lines from different nanowrinkles oriented in the same directions might possess similar polarization profiles (S2-S4, S3-S5), further complicating the process of isolating individual emission lines with polarization filtering. Nevertheless, individual emission lines above 780 nm can be isolated using a long-pass or narrow band-pass optical filter (e.g., S5 in Figure 1d).
To evaluate the single-photon purity of these emission lines, we performed the second-order autocorrelation () measurement from a representative emission line ( nm) under continuous-wave (CW) excitation at 455 nm via a Hanbury Brown–Twiss (HBT) setup. In this setup, the emitted photons are coupled to a 50:50 fiber beam splitter, whose output signals are detected by superconducting nanowire single-photon detectors (SNSPDs). The emission line at 807 nm was spectrally isolated using a 750 nm long-pass filter. Figure 1e shows the obtained correlation histogram, exhibiting a clear signature of photon antibunching. The extracted value of is below the threshold, indicating single-photon emission behavior. The reduced purity obtained from the particular emitter can be attributed to emission leakage from nearby SPEs that fall within the diffraction limit of our collection spot. This could be overcome by spatially isolating the nanowrinkles via novel nanostructure designs and subsequently controlling their directionality. The challenges posed by the lack of control of the localization and polarization in SPEs based on small cylindrical nanopillars emphasize the need for an effective strain-engineering approach to overcome these limitations.
II.2 Control of nanowrinkle orientation via novel nanopillar designs
@@
We propose tailored nanostructure designs consisting of polygonal structures with long and sharp tips. As illustrated in Figure 2a, their shape allows strain to be induced by the vertices, facilitating the creation of spatially isolated nanowrinkles. Commanding the directionality of the one-dimensional SPE-hosting nanowrinkles gives rise to an effective control of the polarization of the emitted single photons So et al. (2021); Azzam et al. (2023); Wang et al. (2021). Here, we explore three different geometries, namely, a three-pointed star (TS), a five-pointed star (FS), and a bowtie (BT). The structures are patterned out of negative e-beam resist (HSQ), which is cured into SiO upon e-beam exposure. The thickness of the nanostructures ( nm) was identified as optimal for producing nanowrinkles during the transfer process. Further details of the e-beam lithography and the remaining fabrication process are described in the Methods section.
@@
First, we demonstrate WSe single-photon emitters with a TS nanostructure. Figure 2a shows a 3D schematic representation of the TS nanostructure with a WSe monolayer flake, illustrating the formation of nanowrinkles that follow the directionality of each arm. The surface morphology of a sample hosting a WSe monolayer flake is shown in the atomic force microscope (AFM) image in Figure 2b. The false coloring highlights the fabricated nanowrinkles versus the planar monolayer, with the inset providing a high-resolution SEM image of the same structure for clarity. We examine the morphological profiles of the fabricated nanowrinkles on the different designs in Figures S3,4,5 of the Supporting Information. We notice clear similarities in the nanowrinkles between the different geometries, meaning that a similar amount of strain is applied to the monolayer WSe regardless of the design. More importantly, the nanowrinkles seem to be forming at the vertices, which was one of the primary reasons for tailoring these designs. Moreover, due to the size of the structures, the nanowrinkles are separated by more than 2 µm, allowing us to address each one individually. Nevertheless, the alignment between the arms and the nanowrinkles is imperfect and will be addressed in the following sections.
To investigate how these nanowrinkles modify the optical properties of monolayer WSe, we performed PL imaging under a 470 nm LED excitation in a closed-cycle optical cryostat equipped with a microscope objective (60, NA = 0.82) operating at T = 4 K. The obtained PL image (T = 4 K) taken from the TS sample is presented in Figure 2c, where increased PL emission is observed along the nanowrinkles. This increased PL emission is a sign of the successful fabrication of SPEs. Interestingly, the emission from the top of the TS structure, highlighted by a semi-transparent star, is comparable to that of the planar monolayer. This observation suggests that the mechanical strain applied at the sides of the arms or the top of the structure may not alter the bands significantly. In addition, forming a nanowrinkle seems crucial to induce the confinement effect, ultimately resulting in the generation of SPEs. Moreover, the increased PL signal is not only focused on the tips of the arms but can also be seen across the length of the nanowrinkles. To gain some insights into the effect of the formed nanowrinkles on the polarization of the emitted photons, we performed the polarization-resolved µPL measurement by mounting a linear polarizer to the characterization setup presented in the Methods section. The collected PL spectra from the individual nanowrinkles (P1-P3) under parallel and orthogonal polarizations with respect to the directionality of each nanowrinkle are presented in Figures 2d-f. All three emitters exhibit highly polarized emission, as the PL intensity of the orthogonal polarization is orders of magnitude lower than the parallel one. To gain a better understanding, we analyzed the polarization-resolved µPL spectra taken from the individual nanowrinkles. The extracted PL intensity as a function of the in-plane polarization exhibits highly linear polarized behavior, as shown in Figure 2g-i. To quantify the degree of linear polarization (DOLP) of the emitted photons, we use a sinusoidal fitting function and extract the DOLP using the maximum and minimum PL intensities as , with the calculated values of 99 ± 4%, 82 ± 14%, and 93 ± 3%, respectively for the three emitters. The strong degree of linear polarization is attributed to the nanowrinkles, which give rise to a localized one-dimensional strain potential, thus forming strongly aligned dipoles. Tripathi et al. (2018); Drawer et al. (2023) In addition, the fitting reveals that the in-plane polarization profile of each emitter does follow the directionality of their respective nanowrinkles. Hence, the proposed design facilitates the successful fabrication of spatially isolated SPEs in monolayer WSe. Despite the effective fabrication of SPEs, the TS structures produce nanowrinkles that are not perfectly aligned with their respective arms. With the purpose of reducing the extension of the nanowrinkles to create spatially confined emitters, we introduce the FS structure. We speculate that by increasing the number of vertices while maintaining the size of the structures, the nanowrinkles will shorten and be more localized around the tips.
@@
In Figure 3, we present similar data for investigating the proposed FS design. The AFM image in Figure 3a showcases the surface morphology of the FS sample and the resulting nanowrinkles. Once more, the false coloring highlights the nanowrinkles, while the false-shading representation is shown in the inset for better contrast. Here, one of the five vertices has no visible nanowrinkle forming at its end. This is considered a normal occurrence due to the intricacies of the transfer process. Moreover, we observe that in the other four vertices (P1-P4), the nanowrinkles are indeed significantly shorter than those in the TS, leading to more localized SPEs. Despite this, the nanowrinkles are still well separated and can be addressed individually.
@@
To further analyze the optical properties of the monolayer WSe on the FS structure, we obtain the corresponding PL image (T = 4 K) under a 470 nm LED excitation, as shown in Figure 3b. The PL signal significantly increases around the tips where the nanowrinkles are located (P1-P4), with the tip missing a nanowrinkle showing emission comparable to the pristine monolayer.
@@
To understand the effect of the nanowrinkles of the FS structure on the in-plane polarization of the single photons and compare it to that of the TS, we obtain polarization-resolved µPL spectra from the four highlighted spots. We extract the PL intensities and present them as polar plots in Figures 3c-f. Similarly to Figure 2, the sinusoidal fitting reveals that the emitted photons from each SPE demonstrate high DOLP, with the extracted values at 86 , 81 , 76 , and 85 , respectively. More importantly, the results reaffirm that the in-plane polarization’s orientation follows the nanowrinkles’ directionality, although our control over this directionality is still unsatisfactory.
In general, the capability of the FS structure to host more localized and highly polarized emitters comes with the trade-off that certain vertices may not be able to form nanowrinkles.
@@
While both the TS and FS structures demonstrate some promising amount of control over both the SPEs’ localization and polarization, it is worth restating that the control of nanowrinkle directionality is not flawless. As shown in Figures 2 and 3, possible misalignments between the arm and the nanowrinkle are possible. This behavior is attributed to the hexagonal lattice structure of the material and how its intrinsic “armchair” and “zigzag” directions Le and Rahman (2013) might affect the final directionality of the nanowrinkles. Moreover, a statistical analysis of the deviation in alignment between the directionality of the arms and the nanowrinkles of the TS and FS structures is presented in Figure S6 of the Supporting Information. The Gaussian fitting of the analysis reveals a central position of . However, the standard deviation of highlights the demand for an improved design. Such structure (BT) is presented in Figure 4. The BT design facilitates the formation of the nanowrinkle between the two triangles, localizing the emitters to the best of our ability. Additionally, any potential misalignment between the vertices and the nanowrinkle would be minimized while isolating the SPEs from any naturally forming deformation in the surrounding area. Figure 4a shows the AFM image of an exemplary BT structure with a monolayer WSe flake. A nanowrinkle connecting the two triangles is clearly visible in both the AFM image and the SEM inset.